OpenStudy (anonymous):

write this equation in standard form.3x squared=3x+6. show your work please :)

jimthompson5910 (jim_thompson5910):

\[\Large 3x^2 = 3x+6\] \[\Large 3x^2 - 3x = 6\] \[\Large 3x^2 - 3x - 6 = 0\]

OpenStudy (anonymous):

Standard form: 3x^2 - 3x = 6

Directrix (directrix):

The standardized form of a quadratic equation is ax^2 + bx + c = 0, where "a" does not equal 0. 3x^2 - 3x - 6 = 0 ---> http://jwilson.coe.uga.edu/emt668/EMAT6680.Folders/Barron/unit/Lesson%206/6.html